OpenStudy (anonymous):

Select the expressions that are equivalent to 10x^0. Select all that apply. 10 0 (10x)^0 10(10x)^0

Directrix (directrix):

Do you have any idea which one it might be? I haven't evaluated them yet.

OpenStudy (anonymous):

0 and (10x)^0

Directrix (directrix):

Will you check and see if you posted the problem correctly? Is there supposed to be a value of x given by which to test these options?

OpenStudy (anonymous):

My bad I had typoed the equation it is 10x^0

Directrix (directrix):

Because, as it stands, 0 is not the same as 10x^2 unless x = 0.

OpenStudy (anonymous):

The equations is 10x^0

Directrix (directrix):

10x^0 = 10 times x to the 0 power. x to the 0 power is 1. So, 10x^0 = 10 which is not the same as 10 x^2. Please check to see that you posted all the options correctly. @BradSuga

OpenStudy (anonymous):

All the options are correct I had mistakenly put 10x^2 the correct one is 10x^0

Directrix (directrix):

Oh, the problem changed to 10x^0. Well, we know that equals 10. So, which of the options are equal to 10?

Directrix (directrix):

I see 2 of them. We already know one.

Directrix (directrix):

Surely, this option: 10 What is the other? @BradSuga

OpenStudy (anonymous):

10(10x)^0

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Let's check 10(10x)^0 = 10 * (10x)^0 = 10 * 1 = 10 Yes, that is correct.

Directrix (directrix):

The first and the last, right?

OpenStudy (anonymous):

Yes!

Directrix (directrix):

Good.

OpenStudy (anonymous):

What about this one @Directrix (4a^2b^2)(-6a^7b^-4)

Directrix (directrix):

What is the question?

OpenStudy (anonymous):

Select the expression that is equivalent to (4a^2b^2)(-6a^7b^-4)

Directrix (directrix):

Options?

Directrix (directrix):

Please check to assure that you wrote this correctly: (4a^2b^2)(-6a^7b^-4)

OpenStudy (anonymous):

Select the expression that is equivalent to (4a^2b^2)(-6a^7b^-4) a)-24a^14/b^8 b)-24a^9b^2 c)-24a^14b d-24a^9/b^2

Directrix (directrix):

What you wrote is supposed to look like this. Attached.

Directrix (directrix):

Hey, would you click on the blue DRAW tab and write by hand how the expression in the problem is supposed to look?

OpenStudy (anonymous):

Here is a screenshot of the problem http://prntscr.com/5nefob

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Much clearer. Let me think.

Directrix (directrix):

What is the next step?